OpenStudy (anonymous):

** I NEED HELP WITH THIS APR QUESTION DONT JUDGE MY RESPONSE *** Bridgette opened a store credit card to purchase a business suit for $463. She put the entire purchase on the credit card. Her APR is 14.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months?

OpenStudy (tkhunny):

We already started this one. How far did you get? i = 0.1499 -- Annual Interest Rate j = i/12 -- monthly Interest Rate Month 0 (Time of Purchase): $463.00 Month 1: 463(1+j) - 0.05*463(1+j) = 463(1+j)(1-0.05) = 463*0.95*(1+j) Month 2: 463*0.95*(1+j)*(1+j) - 463*0.95*(1+j)*(1+j)*0.05 = 463*[0.95*(1+j)]^2 I'm starting to see a pattern. You?